# The Proper Motion of Strongly Lensed Binary Neutron Star Mergers in LIGO/Virgo/Kagra can be Constrained by Measuring Doppler Induced Gravitational Wave Dephasing

### v1

@@ 1 Introduction
In this letter, we show that the planned sensitivity of the advanced LVK facilities over the next years is sufficient to detect GW phase shifts from strongly lensed NS-NS GW mergers, which will allow us to directly measure their relative transverse proper motion. In particular, we focus on the planned A+ and proposed A# sensitivites as detailed in e.g. the T2200287v3 public document. This measurement becomes possible due to the low mass of the NS, which results in a relatively large number of GW cycles for phase-shift detection, combined with the possibility for lens magnifications of order (Smith et al., 2023). Together with the additional detectable GW phase shift from a possible acceleration of the NS-NS GW source (e.g. Tamanini et al., 2020; Tiwari et al., 2024; Samsing et al., 2024a; Hendriks et al., 2024a, b), our study hints that constraining the origin of NS-NS mergers through kinematic measurements will become a reality over the next few years with LVK.
@@ 2 Methods
where the tilde denotes the Fourier space representation, and the mode coefficients are known to very high post-Newtonian (PN) orders (Blanchet, 2014). Other variables are listed in Eq. A1. The Fourier phase is given by:
Here the factor accounts for the sky averaged response of an L shaped interferometer to incident GW (Maggiore, 2007; Robson et al., 2019). In this work, we use third order PN waveforms for binaries with no spin.
To efficiently estimate the detectability of dephasing over a large parameter space, we employ the SNR criterion (Kocsis et al., 2011). It states that a waveform perturbation is detectable if the following inequality is satisfied:
@@
where the super script ‘’ denotes the complex conjugate, and is the noise power spectral density of the detector. Here we take the of the advanced LVK detectors, focusing on A+ and A# sensitivity. Values of SNR higher than a few indicate that a given waveform perturbation may be detectable, without accounting for the problem of degeneracies (Zwick et al., 2023). In this letter, we further employ a Bayesian parameter inference pipeline in order to investigate degeneracies and to validate the SNR results, which is detailed in the appendix A. We note that the goal of this pipeline is simply to show the advantages of jointly inferring the parameters of the reference waveform and Doppler shifted waveform. It is not intended as an exhaustive suite of tests including the effect of varying priors, waveform models, or the details of the dephasing prescriptions, which we leave to future work.
@@ 4 Conclusions
Strongly lensed NS-NS mergers are expected to be detected in the first years after the LVK facilities reach A+ sensitivity. We have shown that, for a significant fraction of the expected source parameter phase space, such signals will showcase detectable GW dephasing signatures due to the peculiar transverse velocity of the lens and source. The prospects are even more promising when considering the proposed A# sensitivity curve. We conclude that the first detection of GW dephasing and the first use of GW measurements to directly infer information of the source’s environment is likely only a few years away, and that further investigating the synergies between GW dephasing measurements and multi-messenger observations in lensed NS-NS mergers will have high scientific yield.

### v2

@@ 1 Introduction
In this letter, we show that the planned sensitivity of the advanced LVK facilities over the next years is sufficient to detect GW phase shifts from strongly lensed NS-NS GW mergers, which will allow us to directly measure their relative transverse proper motion. In particular, we focus on the planned A+ and proposed A# sensitivites as detailed in e.g. the T2200287v3 public document. This measurement becomes possible due to the low mass of the NS, which results in a relatively large number of GW cycles for phase-shift detection, combined with the possibility for lens magnifications of order (Smith et al., 2023). Together with the additional detectable GW phase shift from a possible acceleration of the NS-NS GW source (e.g. Tamanini et al., 2020; Tiwari et al., 2024; Samsing et al., 2024a; Hendriks et al., 2024a, b), our study hints that detections of GW dephasing and GW based constraints on the peculiar velocity of NS-NS merger hosts will become a reality over the next few years.
@@ 2 Methods
where the tilde denotes the Fourier space representation, and the mode coefficients are known to very high post-Newtonian (PN) orders (Blanchet, 2014) and the factor accounts for the sky averaged response of an L shaped interferometer to incident GW (Maggiore, 2007; Robson et al., 2019). Other variables are listed in Eq. A1. The Fourier phase is given by:
In this work, we use third order PN waveforms for binaries with no spin. To efficiently estimate the detectability of dephasing over a large parameter space, we employ the SNR criterion (Kocsis et al., 2011). It states that a waveform perturbation is detectable if the following inequality is satisfied:
@@
where the super script ‘’ denotes the complex conjugate, and is the noise power spectral density of the detector. Here we take the of the advanced LVK detectors, focusing on A+ (Cahillane & Mansell, 2022; Capote et al., 2025) and A# (Gupta et al., 2024) sensitivity. Values of SNR higher than a few indicate that a given waveform perturbation may be detectable, without accounting for the problem of degeneracies (Zwick et al., 2023). In this letter, we further employ a Bayesian parameter inference pipeline in order to investigate degeneracies and to validate the SNR results, which is detailed in the appendix A. We note that the goal of this pipeline is simply to show the advantages of jointly inferring the parameters of the reference waveform and Doppler shifted waveform. It is not intended as an exhaustive suite of tests including the effect of varying priors, waveform models, or the details of the dephasing prescriptions, which we leave to future work.
@@ 4 Conclusions
Strongly lensed NS-NS mergers are expected to be detected in the first years after the LVK facilities reach A+ sensitivity. We have shown that, for a significant fraction of the expected source parameter phase space, such signals will showcase detectable GW dephasing signatures due to the peculiar transverse velocity of the lens and source. The prospects are even more promising when considering the proposed A# sensitivity curve. The methodology proposed here allows for a "dark", or GW based, constraint of the host galaxy transverse velocity within its galaxy cluster. Additionally, not modelling the effects here presented can lead to the misidentification of a lensed GW signal as two distinct sources with different chirp mass, due to the scaling of the Doppler induced dephasing. We conclude that the first detection of GW dephasing and the first use of GW measurements to directly infer information of the source’s environment is likely only a few years away. In light of this, We suggest that further investigating the synergies between GW dephasing measurements and multi-messenger observations in lensed NS-NS mergers should be prioritised.